A scouting troop has been invited to sell ice cream sandwiches as a fundraiser at the city park July 4th fireworks celebration. The sandwiches will cost one dollar each and will be sold for two dollars each. There are no other costs because the city park has agreed to provide use of facilities and freezers. The scouts will have to dispose of, eat, or give away any ice cream sandwiches that are left over after the event. The following "Payout Table" provides values for the events and alternatives of the ice cream venture.

Value of Events Profits in Dollars
Alternatives Events
Ice Cream Sandwich Inventory Ice Cream Sandwich Demand Expected value of Alternative
1,000 (10%) 1,100 (20%) 1,200 (40%) 1,300 (20%) 1,400 (10%)
1,000 $1,000 $1,000 $1,000 $1,000 $1,000 $1,000
1,100 $900 $1,100 $1,200 $1,100 $1,100 A
1,200 $800 $1,000 $1,200 $1,200 $1,200 B
1,300 $700 $900 $1,200 $1,300 $1,300 C
1,400 $600 $800 $1,000 $1,200 $1,400 D

What is the missing value for D, which has an expected value of an inventory of $1,400?

A. $1,000

B. $1,080

C. $1,200

D. $1,120
Answer #1

D = 2*(1000*0.1 + 1100*0.2 + 1200*0.4 + 1300*0.2 + 1400*0.1) - 1400

= $1000

C = 2*(1000*0.1 + 1100*0.2 + 1200*0.4 + 1300*0.3) - 1300

= $1080

B = 2*(1000*0.1 + 1100*0.2 + 1200*0.7) - 1200

= $1120

A = 2*(1000*0.1 + 1100*0.9) - 1100

= $1080